OpenStudy (petlovers123):

Daisy is making an abstract painting with two triangles. The dimensions of the painting are shown below: Two triangles share the same height of 10 inches. One triangle has a base of 8 inches, and the other is 12 inches. The total area of the two triangles is _____ square inches.

OpenStudy (petlovers123):

@Liam2004

OpenStudy (petlovers123):

@mathstudent55

OpenStudy (petlovers123):

@caozeyuan

OpenStudy (caozeyuan):

A=1/2*b*h

OpenStudy (caozeyuan):

since they have same h, we can add b together

OpenStudy (caozeyuan):

so new b=20, 20*10*1/2=100

OpenStudy (misssmartiez):

1st triangle: 8 x 10 = 80 x 0.5 = 40 2nd triangle: 12 x 10 = 120 x 0.5 = 60. 60 + 40 = 100. The total area of the two triangles is 100 square inches.

OpenStudy (misssmartiez):

Using the formula: B * H * 0.5 (OR 1/2)

OpenStudy (misssmartiez):



OpenStudy (misssmartiez):

Get it @petlovers123?

OpenStudy (kevin8767):

64 is your answer if you need explaining please don't hesitate to ask.

OpenStudy (misssmartiez):

explain how.